# Large Language Model Unlearning for Source Code

### v1

@@ I Introduction
  • Extensive experimental results demonstrate that PROD significantly outperforms existing LLM unlearning approaches, achieving successful forgetting of specific code while effectively preserving code generation capabilities.
Our source code and datasets will be released at https://github.com/jiangxxxue/PROD.
II Related Work
II-A Code Generation with LLMs
Code generation is the computational process of automatically producing executable source code from specifications or natural language inputs. Generative models such as ChatGPT [13] and Claude [14] have demonstrated strong capabilities in code generation, benefiting from pretraining on large-scale data and code repositories. Given the immense potential of LLMs, researchers have developed code-specialized LLMs, such as CodeX [15], CodeLlama [16], DeepSeek-Coder [17], Qwen2.5 Coder [18], and CodeGemma [19], which further enhanced LLMs’ capabilities across various software development tasks. However, several critical challenges remain unresolved in leveraging LLMs for code generation, including concerns about safety, legality, and reliability.
Recently, an emerging concern is that LLM-generated code may inadvertently violate security best practices. Because LLMs learn from massive public code repositories, many of which include buggy or outdated patterns, they can propagate known vulnerabilities or poor coding habits. Researchers have begun exploring security-focused code generation techniques [5, 20, 21, 22, 23, 24]. Representative approaches include SaferCode [5], which augments LLM training with carefully curated secure code datasets, and work [20], which proposes incorporating security examples directly into prompts to avoid unsafe coding patterns. Beyond security issues, VersiCode [21] introduces a version-specific benchmark to evaluate how models handle API changes across different releases, revealing that version-controllable code generation poses significant challenges. Failing to adapt can lead to functional errors and long-term maintenance issues. Similarly, SecuCoGen [22] systematically analyzes deprecated API usage in LLM-based code completion. Testing seven advanced models on over 28k prompts and 145 API changes, they observe a substantial rate of outdated suggestions, indicating that many LLMs fail to migrate to newer APIs. Additionally, copyright concerns have also been raised regarding LLM-generated code. LiCoEval [23] systematically evaluates LLMs’ license compliance and finds that models sometimes produce code fragments that closely resemble copyrighted implementations, often without proper attribution.
II-B LLM Unlearning
Machine unlearning, the process of removing specific data and its influence from a trained machine learning model, has garnered significant attention in recent years. Since [25] first introduced the concept of machine unlearning, laying the groundwork for subsequent research. Due to the widespread application and the relative simplicity of classification models, they become the focal point for most of the machine unlearning works [26, 27, 28, 29, 30, 31]. Machine unlearning approaches can be broadly categorized into three types: data-reversed training [32, 33, 34], influence function based approaches [28], and optimization based unlearning [35, 36]. Data-reversed training techniques employ noise injection to reverse or weaken the model’s memory and dependence on specific data. Influence function-based approaches evaluate the impact of training samples on the model by perturbing model parameters, which are less common in LLM unlearning due to the computational complexity of Hessian matrix inversion. Optimization-based unlearning, which is the focus of our work, achieves forgetting through direct model optimization. The rapid advancement of LLMs presents more significant challenges for machine unlearning. Their massive parameter scales and generative nature make traditional unlearning approaches unsuitable [8]. Recent research efforts have been dedicated to the domain of LLM unlearning within the field of NLP, proposing methods such as GA [9], DPO [11], NPO [10], and FLAT [12], which aim to address forgetting tasks in the field of NLP, such as user privacy unlearning, copyrighted books unlearning, and harmful content unlearning [9]. To facilitate the evaluation of LLM unlearning, several benchmarks have been proposed. These include synthetic profiles and question-answer pairs about fictitious authors [37], as well as the forgetting of specific content like Harry Potter books [38]. Moreover, to assess the vulnerabilities of unlearned models, various attack approaches have been employed, such as prefix injections [39, 40, 41]. These attacks help researchers understand the robustness of unlearning techniques and identify areas for improvement.
III Preliminary Investigation on LLM Unlearning
We first give the definition of LLM unlearning problem. Then we introduce four representative LLM unlearning approaches. Finally, we investigate the performance of existing LLM unlearning approaches on code-related tasks and elaborate on the severe degradation issues in model utility caused by existing unlearning approaches when applied to source code.
III-A Problem Definition
For the unlearning problem, we begin with forget data $D_{f}$ and a LLM $\pi_{\theta}$ that needs to undergo unlearning. $D_{f}$ represents the content we wish the model to forget. In practical scenarios, $D_{f}$ might be a collection of model-generated code that has been reported by users for copyright infringement. $D_{f}$ consists of inputs $x_{f}$ and outputs $y_{f}$, where the inputs represent prompts that might trigger undesired outputs $y_{f}$ [^2] Inputs may be omitted in some scenarios, such as when the goal is to forget entire code content, the prompt can be empty.. $D_{f}$ does not need to be an exact subset of $\pi_{\theta}$’s training data. The objective of LLM unlearning is to efficiently remove the impact of $D_{f}$ on $\pi_{\theta}$, meaning that the unlearned model $\pi_{\theta}^{*}$’s outputs should be as distant as possible from $y_{f}$.
We evaluate the effectiveness of an LLM unlearning approach from two dimensions: forget quality and model utility. Forget Quality refers to the degree to which the unlearned model eliminates the influence of target forget data, while Model Utility refers to the general performance of the model on other tasks. An ideal unlearning approach aims to increase forget quality while minimizing the reduction in model utility.
III-B LLM Unlearning Approaches

### v2

@@ Introduction
  • Extensive experimental results demonstrate that PROD significantly outperforms existing LLM unlearning approaches, achieving successful forgetting of specific code while effectively preserving code generation capabilities. Our source code and data are available at https://github.com/jiangxxxue/PROD.
Related Work
In this section, we outline the most relevant work of PROD and detail the extended related work section in Appendix.
Code Generation with LLMs
Code generation has been significantly advanced by LLMs, from general-purpose LLMs like GPT-4 (achiam2023gpt4) to a variety of specialized code LLMs (codex; CodeLlama; zhu2024deepseekcoder; hui2024qwen2.5-coder; team2024codegemma). However, their training on vast, public code repositories introduces critical challenges regarding security, reliability, and legality. A primary concern is that LLMs may replicate vulnerabilities. To mitigate this, researchers have explored security-focused fine-tuning with curated datasets (SaferCode (safecoder)) and in-context learning with secure examples (mohsin2024can). Beyond security, model reliability is undermined by poor handling of software versions. Studies show that LLMs often suggest deprecated APIs (SecuCoGen (wang2024and)) and struggle with API changes across different library releases, a challenge highlighted by the VersiCode benchmark (wu2024versicode). Finally, legal issues arise from models generating code that may violate software licenses, as systematically evaluated by LiCoEval (xu2024first).
These works highlight that despite remarkable progress, LLM-based code generation still falls short of developers’ expectations for producing secure, legally compliant, and up-to-date code. Our work aims to address this gap by proposing a general unlearning approach for suppressing undesired code output from LLMs.
LLM Unlearning
Machine unlearning aims to remove the influence of specific data from a trained model (DBLP:conf/sp/CaoY15). Initial research primarily focused on classification models (DBLP:conf/cvpr/GolatkarAS20; DBLP:conf/aistats/IzzoSCZ21; DBLP:conf/sp/BourtouleCCJTZL21), with methods broadly categorized into data-reversed training (DBLP:journals/tnn/TarunCMK24; DBLP:journals/tifs/ChundawatTMK23), influence function-based approaches (DBLP:conf/aistats/IzzoSCZ21), and optimization-based unlearning (DBLP:conf/icml/GuoGHM20; DBLP:conf/alt/Neel0S21).
The vast scale and generative nature of LLMs present unique challenges, rendering many traditional methods impractical (DBLP:journals/natmi/LiuYJCBHYLXLVBKL25). Specifically, the computational cost of Hessian inversion makes influence function-based methods prohibitive for LLMs. Consequently, recent work has shifted towards optimization-based techniques tailored for LLMs in the NLP domain. Methods such as GA (Unlearning), DPO-based unlearning (DPO), NPO (NPO), and FLAT (FLAT) have been proposed to erase private data, copyrighted material, and harmful content. The evaluation of these techniques relies on specialized benchmarks (DBLP:journals/corr/abs-2401-06121; DBLP:journals/corr/abs-2310-10683) and is often tested for robustness against adversarial attacks like prefix injections (Jailbroken; DBLP:journals/corr/abs-2408-10682; qi2025safety).
Given that source code possesses unique characteristics compared to natural language, although the exploration of LLM unlearning in NLP has proven its value, its application in code generation remains largely unexplored.
Preliminary of LLM Unlearning
In this section, we first formalize the LLM unlearning problem and establish the evaluation criteria. Next, we briefly survey four representative baselines that span current LLM unlearning paradigms. Finally, we conduct an empirical pilot study revealing that, on source code, existing baselines cause great losses in model utility, motivating the need for a more surgical solution.
Objective of Unlearning
Within the broad landscape of machine unlearning, we focus on optimization-based methods. Here, unlearning is a post-training procedure that removes the LLM’s ability to produce specific, undesirable code snippets. Let $\mathcal{D}_{f}=\{x_{f}^{(i)},y_{f}^{(i)}\}_{i=1}^{N_{f}}$, denote the set of such snippets, where each pair consists of a prompt $x_{f}^{(i)}$, and its undesirable completion $y_{f}^{(i)}$ (note that the prompt $x_{f}^{(i)}$ can be empty in cases such as the entire snippet $y_{f}^{(i)}$ must be forgotten). After unlearning, the target LLM $\pi_{\Theta}$ is supposed to assign zero probability to every undesirable $y_{f}^{(i)}$ given $x_{f}^{(i)}$. Formally, we quantify this objective with the loss $\mathcal{L}_{f}\left(\pi_{\Theta}(x_{f}),y_{f}\right)$ that measures the divergence between the LLM’s output $\pi_{\Theta}(x_{f}^{(i)})$ and the undesirable output $y_{f}^{(i)}$. Unlearning is then cast as the following optimization:
$\min_{\Theta}\sum_{i=1}^{N_{f}}-\mathcal{L}_{f}\left(\pi_{\Theta}(x_{f}^{(i)}),y_{f}^{(i)}\right),$ (1)
where maximizing $\mathcal{L}_{f}$ forces $\pi_{\Theta}$ to forget every undesirable $y_{f}^{(i)}$ in $\mathcal{D}_{f}$, including copyrighted code, insecure fragment, or deprecated API invocation, etc. Notably, $\mathcal{D}_{f}$ merely captures behaviors we wish to eliminate, and samples in it do not have to appear verbatim in the original training data.
Metrics
As aforementioned, unlearning requires to erase only what is undesirable or even harmful, and retain everything else. We therefore assess unlearning approaches along two axes: ❶ Forget quality: the thoroughness with which $\pi_{\Theta}$ eliminates the influence of $\mathcal{D}_{f}$, measured by the degree of dissimilarity between the output of $\pi_{\Theta}$ and $y_{f}^{(i)}$ when prompted with $x_{f}^{(i)}$. ❷ Model utility: the collateral side damage caused by the procedure, gauged by $\pi_{\Theta}$’s retained performance on other general tasks and datasets. Assessing the balance between forget quality and model utility is essential due to their trade-off in unlearning methods.
Figure 1: A case on the forget quality and model utility of existing method in unlearning code. Existing methods (GA, NPO, FLAT, DPO) exhibit severe utility degradation, such as mute refusal, token collapse, or syntactic incoherence. PROD successfully unlearns copyrighted content while maintaining utility for general code generation tasks.
LLM Unlearning Solutions
Existing methods are almost exclusively tuned to optimize the forget objective in Equation (1), and treat downstream utility as an after-thought, resulting in plausible damage. ❶ Gradient Ascent (GA) (Unlearning) carries out a “reverse” of training by ascending the loss on $\mathcal{D}_{f}$, pushing $\pi_{\Theta}$ away from the parameter region that supports undesirable generations. ❷ Negative Preference Optimization (NPO) (NPO), which is a variant of Direct Preference Optimization (DPO) (DPO), supplies only negative examples (i.e., $(x_{f}^{(i)},y_{f}^{(i)})$) to repel $\pi_{\Theta}$ from the unwanted distribution. ❸ DPO is a general preference optimization method that, while not specifically designed for unlearning, can be adapted for this purpose by treating the content-to-be-forgotten as the “dispreferred” response. ❹ Forget Data Only Loss Adjustment (FLAT) (FLAT) formulates $\mathcal{L}_{f}$ with $f$-divergence, enforcing $\pi_{\Theta}$ to assign maximal probability to a hand-crafted refusal template (e.g., “I do not know”) while minimizing probability on each $(x_{f}^{(i)},y_{f}^{(i)})$.
Figure 2: The PROD unlearning pipeline. The pipeline contains three key steps: ❶ Suppress the probabilities of undesirable code snippets to zero. ❷ Redistribute the probabilities across the remaining vocabulary. ❸ Optimize the model to match the surgically sculpted target distribution.
A Motivating Example
To vividly show the collateral damage caused by existing unlearning methods on source code tasks, we conduct an empirical study. We benchmark the above-mentioned baselines on a copyrighted code unlearning task and then measure their retained utility with a standard code generation benchmark (detailed setups in the experiment section). As Figure 3 indicates, all baseline approaches achieve $\approx 90\%$ forgetting ratio, yet their success ratio on code generation degrades to near-zero when they reach best forget quality. In other words, these approaches successfully erase the undesirable snippets (i.e., copyrighted code) but simultaneously cause LLMs to forget how to program, i.e., such an outcome makes the target LLMs almost useless. Our motivation in this paper is precisely the opposite, as we argue to excise the undesirable code snippets while leaving the target LLMs fully operational.
[…]
PROD
To protect model utility while guaranteeing forgetting, we introduce PROD, a surgically precise unlearning method that manipulates the LLM’s token-level output distributions. As illustrated in Figure 2 and formalized in Algorithm 1, the working pipeline of PROD consists of three steps. ❶ Suppress the probabilities predicted by the target LLM of tokens that constitute the undesirable snippet to zero (i.e., remove them from the output). ❷ Redistribute the probabilities across the remaining vocabulary so that the resulting distribution mirrors the original statistical patterns and preserves knowledge that the LLM learn from the code corpora. ❸ Optimize the target LLM to match this surgically sculpted target distribution. As a result, PROD achieves complete erasure of undesirable or even harmful code with virtually no collateral damage to any other programming competency.
Target Distribution Sculpting
We begin by sculpting the target LLM’s output distribution into a precise supervisory signal that will steer the entire unlearning process of PROD.
Output Distribution Collection
The initial step of PROD retrieves the LLM’s raw next-token distributions for each $(x_{f},y_{f})$ pair  [^1] For simplicity, the superscripts in $x_{f}^{(i)}$ and $y_{f}^{(i)}$ are omitted henceforth in this section.. Given a forget pair $(x_{f},y_{f})$ with $y_{f}=(y_{f,1},\cdots,y_{f,L})$, at time-step $t$, we feed the target LLM ($\pi_{\Theta}$) the concatenation of the prompt $x_{f}$ and the preceding ground-truth prefix $y_{f,<t}=(y_{f,1},\cdots,y_{f,t-1})$ to obtain the logits $h_{t}\in\mathbb{R}^{|\mathcal{V}|}$, where $\mathcal{V}$ refers to the vocabulary. A softmax converts the logits into the original distribution $p_{t}(\cdot|x_{f},y_{f,<t})$. $p_{t}$ then serves as the canvas where the subsequent manipulations will be carried out.
Forget Data Elimination
To prohibit any generation of the target sequence, PROD performs a single, decisive edit on $h_{t}$ at every $t$. The logit entry corresponding to the target token $y_{f,t}$ is suppressed to -$\infty$ as below:
$\hat{h}_{t}[j]=\begin{cases}-\infty,&\text{If~}\mathcal{V}[j]=y_{f,t},\\ h_{t}[j],&\text{Otherwise},\end{cases}$ (2)
where $\cdot$] denotes vector indexing, and $\mathcal{V}[j]$ refers to the $j$-th token in the whole vocabulary $\mathcal{V}$. A softmax normalization of $\hat{h}_{t}$ would yield the eliminated distribution $\hat{p}_{t}(\cdot|x_{f},y_{f,<t})$, which assigns zero probability to $y_{f,t}$ and redistributes its entire mass across the remaining vocabulary in exact proportion to their original likelihoods.
Noise Elimination
Suppressing the target token $y_{f,t}$ can inadvertently amplify less probable, non-target tokens, injecting noise into $\hat{p}_{t}(\cdot|x_{f},y_{f,<t})$. We therefore prune the tail (i.e., the noise) via nucleus sampling (Holtzman) with the threshold of $p$. Let $\mathcal{S}_{p}$ denote the smallest set of tokens whose cumulative probability under $\hat{p}_{t}$ reaches $p$. All logits outside $\mathcal{S}_{p}$ are then set to -$\infty$:
$\widetilde{h}_{t}[j]=\begin{cases}\hat{h}_{t}[j],&\text{If~}\mathcal{V}[j]\in\mathcal{S}_{p},\\ -\infty,&\text{Otherwise.}\end{cases}$ (3)
Only tokens within $\mathcal{S}_{p}$ receive redistributed probability, and the rest are eliminated as noise. The softmax normalization upon $\widetilde{h}_{t}$ results in noise-trimmed distribution $\widetilde{p}_{t}$.
Crucially, forget data elimination must precede noise elimination. Reversing the order would either retain target tokens or produce an overly sparse distribution. This two-step sequence ensures a clean unlearning process while preserving a robust, information-rich distribution over the surviving vocabulary.
Probabilistic Redistribution
After excising both the target token $y_{f,t}$ and the noisy tail outside $S_{p}$, $\widetilde{h}_{t}$ holds the remaining probability mass. PROD redistributes this mass across the surviving safe vocabulary (i.e., $\mathcal{S}_{p}$) through softmax, and therefore adopts $\widetilde{p}_{t}$ as the supervisory signal $p_{T}$.
Post-Training
With the sculpted supervisory distribution $p_{T}=\widetilde{p}_{t}$, PROD optimizes the target LLM $\pi_{\Theta}$ via standard cross-entropy loss, nudging its outputs to align with this sculpted distribution while preserving every other learned competency.
Objective
The optimization objective of PROD is to minimize the loss $\mathcal{L}_{\text{{\sc PROD}}}$, finalized as following:
$\mathcal{L}_{\text{{\sc PROD}}}(x_{f},y_{f};\pi_{\Theta})=-\sum_{t=1}^{L}\sum_{w\in\mathcal{V}}p_{T}(w|\cdot)\log\pi_{\Theta}(w|\cdot),$ (4)
where $\pi_{\Theta}(\cdot)$ refers to the actual output distribution of $\pi_{\Theta}$. The conditions (i.e., $x_{f},y_{f,<t}$) are neglected for simplicity. In default design, PROD adopts cross-entropy. The framework is agnostic, i.e., KL (kullback1951), JS (lin1991), or any other distributional divergence can be substituted seamlessly.
Optimization
Any compatible optimizer, with gradient descent included, can be employed in PROD.
Algorithm 1 Pseudocode for PROD.
0: LLM $\pi_{\Theta}$; forget set $\mathcal{D}_{f}$; learning rate $\eta$; training epochs NN; hyperparameters $p,\alpha$.
0: Unlearned LLM $\pi_{\Theta}^{*}$.
1:for epoch k =1=1 to NN do
2:  for each $(x_{f},y_{f})\in\mathcal{D}_{f}$ do
3:   Collect output distribution $p(w|x_{f},y_{f,<t})$.
4:   Compute $\hat{h}_{t}$ to apply forget data elimination.
5:   Compute $\widetilde{h}_{t}$ to perform noise elimination.
6:   Compute probabilistic redistribution $p_{\text{T}}$.
7:   Calculate loss function $\mathcal{L}_{\text{PROD}}$.
8:   $\Theta_{k+1}=\Theta_{k}-\eta\nabla_{\Theta}\mathcal{L}_{\text{PROD}}.$
9:   $k\leftarrow k+1$ and $\Theta^{*}\leftarrow\Theta_{k}$
10:  end for
11:end for
12:return $\pi_{\Theta}^{*}$
Key Know-How of PROD
In general, PROD provides two strengths. ❶ Model utility preservation: the supervisory $p_{T}$ is a surgically sculpted clone of the original $p_{t}$, excising only the undesirable or forbidden snippets while leaving every other competency untouched. ❷ Efficient convergence: since $p_{T}$ remains close to the $\pi_{\Theta}$’s initial output, the optimization process is fast, stable, and light on compute. Furthermore, we elaborate on two implementation details.
$\mathbf{\alpha}$-suppression
Although PROD excels in unlearning long code, early experiments revealed that short forget sequences $y_{f}$ often require more training steps since the supervisory signal is weak. To modulate the “forget strength”, we introduce the $\alpha$-suppression trick and modify the supervisory signal $p_{T}$ as follows:
[…]
where $p_{o}$ and $\widetilde{p}_{t}$ refers to the original and noise-trimmed distributions, respectively. Still, some conditions are neglected for better presentation. $\alpha$ is a hyperparameter to control the degree of forget, i.e., setting $\alpha=0$ disables the trick ($p_{T}=\widetilde{p}_{t}$); increasing $\alpha$ amplifies the negative pressure on $y_{f}$, yielding a stronger unlearning signal. Note that while $\alpha$-suppression introduces negative probability, it is harmless, or even beneficial, under the setting of cross-entropy as defined in Equation 4: ❶ the remaining terms maintain the loss magnitude and gradient direction; and ❷ the negative entry amplifies the penalty on the forget token without computational issues.
Training Stability
As training proceeds, $\pi_{\Theta}$’s distribution drifts with every gradient step. If we recollected $p_{o}$ and recomputed $p_{T}$ on-the-fly, the supervisory signal would swing constantly, destabilizing the training process and risking collapse. To avoid such an issue, we freeze $p_{o}$ at its initial state, i.e., the distributions produced by the target LLM before unlearning, thereby guaranteeing a steady and reliable target throughout optimization.
(a) Copyrighted Code Unlearning
(b) Insecure Code Unlearning
(c) Deprecated API Unlearning
Figure 3: Forget quality versus model utility across different unlearning tasks. Performance curves positioned closer to the upper-right corner indicate superior approach effectiveness.
Evaluation
We introduce a dedicated benchmark for code unlearning and report the main experimental results in this section. Please refer to the appendix for full implementation details and a perceptual quality study.
Code Unlearning Benchmark
Current unlearning benchmarks focus almost exclusively on natural language, leaving code generation unexamined. To fill this gap, we curate and adapt datasets into a code unlearning benchmark, paired with custom metrics. The benchmark contains three unlearning tasks (i.e., copyrighted code, insecure code, and deprecated APIs) to measure forget quality and a general code generation set to assess model utility.
Copyrighted Code Unlearning
We randomly draw $100$ files from the high-quality, deduplicated Stack corpus (TheStack) to serve as the forget set $\mathcal{D}_{f}$, simulating scenarios where users wish to purge copyrighted code files from an LLM. To mimic copyright protection, test files, toy examples, and standard templates are filtered and discarded, retaining those with expressive form, such as substantive implementations, business logic, and non-trivial algorithms. During unlearning, $x_{f}$ is empty and $y_{f}$ contains the full content of the retrieved file; when assessing forget quality, we prompt the target LLM with the first half of each file and evaluate against the other half. As the law of copyright protects expression instead of ideas (autry2002toward; copyright_protect), textual similarity is employed as an indicator of potential copyright infringement. Thus, the forget quality score is measured by the complement of BLEU (BLEU) (i.e., $1-\text{BLEU}$), where higher values indicate stronger erasure of the protected code snippets.
Insecure Code Unlearning
CyberSecEval from Purple Llama (PurpleLlama) is a cybersecurity benchmark, containing 1,916 code snippets collected from open sources that exhibit 50 distinct CWE vulnerabilities (cwe) across eight programming languages. We adopt these snippets as our forget set $\mathcal{D}_{f}$. Evaluation proceeds in an autocomplete setup, i.e., given the code that precedes a known vulnerable segment (i.e., $x_{f}$), we prompt the model to continue and then inspect the continuation for security flaws (i.e., $y_{f}$). Forget quality is assessed by the average of two complementary indicators: ❶ $1-\text{BLEU}$ between the LLM’s generation and the vulnerable snippet, and ❷ the pass rate under CyberSecEval’s built-in static analyzer, which detects any insecure lines.
Deprecated API Unlearning
This task is curated atop Versicode (wu2024versicode), which is a version-aware code-generation dataset. With the original dataset, we apply three additional preprocessing steps. ❶ Version filtration discards packages with ambiguous version constraints (e.g., >=2.1.0), retaining those with explicit version requirements (e.g., ==2.1.0). ❷ File filtration removes packages that have less than 3 snippets. ❸ Temporal split designates a middle release as the deprecation boundary, labeling all invocations of APIs prior to that release as “deprecated” and the remainder as “still-valid.” The final dataset consists of 252 packages, 3,449 snippets in total ($\approx$14 snippets and 10 distinct APIs per package). Unlearning is carried out on the deprecated APIs, and evaluated on the remaining still-valid ones. Given the code context immediately preceding an API call, we measure forget quality by exact-match accuracy.
General Code Generation
Following prior work (safecoder), we estimate model utility through HumanEval (codex), which is a widely used code generation benchmark. We employ functional correctness verified against test cases as the indicator of model utility.
Overall Performance Indicator
Forget quality and model utility often pull in opposite directions; naively averaging them or even independently assessing them overlooks the real trade-off. To make this multi-objective comparison explicit, we introduce the Pareto Dominance Ratio (PDR). Let $\mathcal{M}=\{m_{1},\cdots,m_{n}\}$ be the set of evaluated unlearning methods and $\mathcal{O}=\{\text{forget, utility}\}$ the two objectives. For any method $m_{i}\in\mathcal{M}$, PDR is defined as following:
$\text{PDR}(m_{i})=\frac{\bigl|{m_{j}\in\mathcal{M}\setminus{m_{i}}\mid m_{i}\succ_{\mathcal{O}}m_{j}}\bigr|}{|\mathcal{M}|-1},$ (6)
where $m_{i}\succ_{\mathcal{O}}m_{j}$ means that $m_{i}$ Pareto-dominates $m_{j}$, i.e., $m_{i}$ is no worse on both objectives and strictly better on at least one than $m_{j}$. Geometrically, on a 2-D scatter of forget quality versus model utility, $m_{i}\succ_{\mathcal{O}}m_{j}$ places $m_{i}$ strictly to the upper-right of $m_{j}$. PDR therefore reports the fraction of rival methods that $m_{i}$ dominates, yielding a single, interpretable summary of overall performance.
Table 1: PDR (in $\%$) among unlearning approaches.
Task GA DPO NPO FLAT Ours
Copyright 15.3 3.6 0.6 0.0 41.8
Insecurity 13.3 0.0 24.5 6.1 70.4
Deprecation 31.5 20.6 51.6 0.0 58.0
Experimental Results
Unlearning Performance
We benchmark PROD against four representative baselines (i.e., GA, DPO, NPO, and FLAT) on all three unlearning tasks. The target model is CodeLlama-7B (CodeLlama). Greedy decoding (zero temperature) is applied for every inference run. On each task, we compute PDR with forget quality (task-specific) and overall model utility (via HumanEval). Comprehensive results are presented in Table 1 and illustrated in Figure 3.
PROD consistently occupies the upper-right region of the scatter plot (Figure 3), simultaneously delivering stronger forget quality and higher model utility than the baselines. The PDR metric (Table 1) also quantifies this dominance – PROD tops every task, yielding an average relative gain of 124% over the strongest competitor. Training dynamics offer further insight. By analyzing the training log, we find that PROD forgets rapidly during early steps and then gently converges on both objectives, achieving high forgetting efficiency without triggering model collapse.
The three tasks pose different levels of challenges, as the curve slope in Figure 3 varies with the volume of undesirable content. Copyrighted code unlearning is the most challenging from this point of view, because it requires LLMs to forget entire files. Remarkably, PROD retains a notable advantage even on this most difficult task.
Application on Different LLMs
We extend our evaluation to four widely-used code LLMs (i.e., CodeLlama-7B (CodeLlama), Qwen2.5-Coder-7B (hui2024qwen2.5-coder), Deepseek-coder-6.7B (zhu2024deepseekcoder), and Starcoder-7B (starcoder)) to demonstrate the versatility of PROD across diverse architectures and training recipes. While holding all other experimental settings fixed, we track both forget quality and model utility over successive epochs on the copyrighted-code unlearning task.
Figure 4: Performance of PROD on different LLMs, where the dotted line represents the best model utility of baselines in achieving comparable (i.e., 90%) forget quality.
Figure 4 reveals a consistent pattern across all four LLMs. As training proceeds, each LLM attains near-perfect forget quality while its utility curve stabilizes, which significantly outperforms all baseline approaches in model utility. This uniformity underscores both the strength and robustness of PROD. Among the evaluated models, CodeLlama-7B and StarCoder-7B exhibit the mildest utility degradation, implying that an LLM’s intrinsic stability can modulate the effectiveness of unlearning.
Adversarial Attacks
Recent studies have exposed the brittleness of unlearning: forgotten content can resurface when an adversary manipulates the prompt (DBLP:journals/corr/abs-2408-10682; qi2025safety). We therefore assess PROD against prefix injection attack, following previous work (qi2025safety). In such an attack, an adversary prepends an increasing number of tokens from the copyrighted snippet to the prompt in an attempt to recover the erased snippets. We conduct this evaluation on the copyrighted code task against CodeLlama-7B and measure robustness with the drop in forget quality. To ensure validity of the experiment (i.e., the model remains practically useful), we only evaluate checkpoints that retain at least 60% of the original LLM’s HumanEval performance.
Figure 5 shows that PROD is significantly more robust under prefix injection attack than other baselines. Specifically, the similarity between the generation and the undesirable copyrighted code stays below 0.05, while others exceed 0.3. Moreover, PROD shows the smallest variance among baselines, underscoring its stable resilience against prefix inject attack. These results provide clear evidence that PROD delivers stronger and more reliable forgetting than baselines.
(a) Loss Function
(b) Hyperparameter $\mathbf{p}$
(c) Hyperparameter $\mathbf{\alpha}$
Figure 6: Ablation results on alternative loss function, and the impact of hyperparameters $p$ and $\alpha$ in PROD.
Ablation Study
Finally, we perform some ablations. We replace the default cross-entropy loss with two alternative divergence measures (i.e., KL and JS). We also systematically vary the key hyperparameters $p$, which governs the strength of noise elimination, and $\alpha$, which scales the amplified suppression on forget data samples.
From Figure 6(a), we observe that Cross-entropy outperforms both KL and JS divergence, delivering the best trade-off between forget quality and model utility. Figure 6(b) shows that the moderate noise elimination ($p=0.8$) achieves optimal results, while no noise elimination yields the worst. Figure 6(c) shows that $\alpha$ has a large impact on unlearning performance. When $\alpha=0$, which prohibits the generation of forget data by setting its probability to zero in the target distribution, the model achieves the most balanced performance. Positive $\alpha$ values are beneficial for deepening the degree of forgetting, while small negative $\alpha$ values delay forgetting to maintain usability.
Conclusion
In this paper, we have investigated existing LLM unlearning approaches on source code, identifying a significant utility degradation problem. To this end, we propose PROD, a surgical, code-oriented unlearning method that forgets targeted code snippets while preserving other knowledge of programming languages intact. We also establish a benchmark covering copyrighted code, insecure code, and deprecated API unlearning tasks. Extensive experiments show PROD significantly outperforms existing methods in both forgetting performance and user experience, while also exhibiting broad applicability and robustness against unlearning attacks.
Acknowledgments
This research is supported by the National Key R&D Program under Grant No. 2023YFB4503801, the National Natural Science Foundation of China under Grant No. 62192733, 62192730, 62192731, and the Major Program (JD) of Hubei Province (No.2023BAA024).
Appendix A Preliminaries on LLM Unlearning Approaches